OpenStudy (anonymous):

MEDALS AND FANS WILL BE GIVEN!!!! The measure of a supplement of an angle is 35 more than twice the complement of the angle. Find the measures of the angle, its supplement, and its complement.

OpenStudy (anonymous):

180-x = 35 + 2(90-x) solve for x

OpenStudy (anonymous):

thats what i did but i ended up with 35-x=-x which makes no sense

OpenStudy (anonymous):

you forgot to distribute the 2 on the right hand side 180-x = 35 + 180 - 2x

OpenStudy (anonymous):

i got 35-x=0?

OpenStudy (anonymous):

x = 35

OpenStudy (anonymous):

oh is that x=35

OpenStudy (anonymous):

kk thanks!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Don't forget the supplement to that angle

OpenStudy (anonymous):

and the compliment

OpenStudy (anonymous):

yeah i got it thanks!!!!